You are here: Foswiki>PartnerObserving Web>ItalyQueue2013>ObsInaf20130406 (15 Apr 2013, EleanoraSani)Edit Attach

INAF queue observing, 2013 April 4 - April 5

Observer: E. Sani, M. Faccini, A. Bongiorno
Telescope Operator: D. Gonzales Huerta
Support Astronomer: D. Thompson
Instrument: LBC

Summary

At sunset we have clear sky and wind gusts between 15 m/s and 20 m/s.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Nascimbeni_DDT HAT-P-12 U-Special F972N20 2h43m30s blue 2h55m30s red COMPLETED
Grazian30A GOODS-North U-scpec r-Sloan 30m  

Details

01:21 during the sunset gusts over 22m/s don't allow us to open the dome
  • we run 25Bias_Bino.ob

04:09 the TO opens the DOME

4:11 Slewing to Nascimbeni_DDT
  • dofpia, /first, /X2
  • seeing from the DIM=1.4
  • Run 00_POINTING.ob
  • lbcrangebal B=042305 R=042300

04:23 Run DEFOCUS.ob to find the right focus
  • ref3 in the blue channel (66000) and the target in the red channel (66000) have >35000 counts
  • we apply -0.2 of defocusing
  • Run DEFOCUS.ob
  • ERROR: no fits files written in /newdata/
  • Run again DEFOCUS.ob with -0.2 defocusing
  • blue channel is ok while the target in the red channel is still >35000 counts
  • we apply another -0.2 of defocusing only in the red channel
  • the red channel is still >35000 counts
  • we apply another -0.1 of defocusing only in the red channel
  • both channels are ok. counts < 35000 as required by the PI
  • The final defocusing are -0.2 in the blue and -0.5 in the red

04:58 we undo the -0.2 in the blue and the -0.5 in the red to redo the dofpia because of coma aberration present in field

04:59 dofpia, /X2

05:05 defocus -0.2 in the blue and -0.5 in the red

05:06 run 01_OBJECT.ob
  • guide star deactivated

05:36 stopped as required by the PI

05:36 run 02_OBJECT.ob
  • cirrus passing by

06:06 stopped as required by the PI

06:06 run 03_OBJECT.ob
  • frame 061308 in the blue channel has 40000 counts in the third ref star
  • frame 061511 in the blue channel has 43000 counts in the third ref star
  • we defocus the blue channel of an additional -0.1 (now we have -0.3 in the blue and -0.5 in the red)
  • frame 061716 is now ok in all 4 ref

06:36 stopped as required by the PI

06:36 run 04_OBJECT.ob

07:06 stopped as required by the PI

07:06 run 05_OBJECT.ob
  • frame 070826 in the blue has 36000 counts in the third ref star
  • frame 071031 in the blue is ok
  • thick clouds passing during the last frames

07:36 stopped as required by the PI

07:36 run 06_OBJECT.ob
  • 073828- blue and 073808- red objects almost not visible because of clouds
  • thick clouds passing during the whole exposure

08:06 stopped as required by the PI

08:06 run 07_OBJECT.ob
  • thick clouds passing during the whole exposure
  • 081425 blue are 081604 red: first with stars visible again

08:36 stopped as required by the PI

08:36 run 08_OBJECT.ob
  • counts in the third ref star in the blue 47000
  • we apply a further defocusing of -0.1 (now we have -0.4 in the blue and -0.5 in the red)
  • counts in the third ref star in the blue are still too high
  • we apply a further defocusing of -0.1 (now we have -0.5 in the blue and -0.5 in the red)

09:11 stopped because of end of transit

09:16 dofpia, /X2 on this scientific field (since here we see some stars - still cloudy sky)
  • dofpia crashed

09:19 dofpia, /X2 on this scientific field (since here we see some stars - still cloudy sky)
  • dofpia crashed

09:22 Slewing to Grazian field
  • preset canceled by the TO to unwrap the azimut

09:29 Slewing to Grazian field
  • dofpia, /X2 on the scientific field of Grazian
  • dofpia crasched
  • dofpia, /X2 on the scientific field of Grazian
  • dofpia crasched
  • TO points the telescope to a field close by where there are bright stars to try to collimate (bright star WT10_385)
  • dofpia, /X2 on this field converged!

09:48 Slewing to Grazian field
  • dofpia, /X2 on the field of Grazian. converged!
  • lbcrangebal B=094851 R=094846

09:52 RUN grazian_GOODSN_UR_27.ob
  • The seeing at the beginning is ~0.9 in red and ~1 in the blue channel
  • thin sky, partly clear
  • The seeing at the end is ~1 in both channels

10:38 We ask the TO to switch to MODS.
  • this is due to clouds that do not allow to obtain images within the specs.

10:52 The TO closes the dome because of rain risk

11:40 TO declares finished the night

-- EleanoraSani - 05 Apr 2013
Topic revision: r6 - 15 Apr 2013, EleanoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback